Solve the equation log^9(3x+14)-log^95=log^92x.
Answer by venugopalramana(3286)

log^9(3x+14)-log^95=log^92x.
i think you mean all logs are to base 9...in that case since all are to a common base we can drop the base.
log(3x+1)-log 5=log(2x)....or.....log(3x+1)-log 5-log(2x)=0
log{(3x+1)/(5*2x)=0..taking antilogs,we get
(3x+1)/10x=1
3x+1=10x...or...10x-3x=1...
7x=1
x=1/7